# Cosmological simulation of radio synchrotron bridge between pre-merging galaxy clusters

### 10 Mar 2026 v1

@@ 3.2.2 Spectral index map
We find that both the observed and simulated distributions have peaks around $\alpha_{60}^{140}\approx-1.5$. The simulated distribution is shifted to slightly larger $\alpha_{60}^{140}$ when the observational sensitivity is taken into account. The observed distribution is more extended toward the large-$\alpha_{60}^{140}$ (flat-spectrum) than the simulated one [^2] Note that the observational error is relatively large for the flat-spectrum points ($\alpha_{60}^{140}>-1$) compared to steeper spectrum ones. See Fig. 2 of Pignataro et al. (2024a).. One of the possible causes for this lack of flat spectra would be our simplified assumption on the CR injection. Although we only consider an aged population of CREs injected at $z=1$, there should also be CREs freshly accelerated at shocks and AGNs at later times. The mixing of continuous injection and reacceleration is expected to generate different components with a broader spectral distribution and a stretched spectrum (e.g., Brunetti et al. 2017; Nishiwaki et al. 2021). In addition, our magnetic-field dynamo model may underestimate the dispersion of the magnetic field strength in the bridge region compared to the resolved dynamo in the MHD simulation (see Fig. 2). A larger field fluctuation may lead to a broader distribution of the spectral index.
@@ 3.3 Bridge profiles
The right panels of Fig. 8 show the profiles perpendicular to the projected merger axis. We consider a region with a width of 2 Mpc centered on the merger axis and a length of 1.2 Mpc parallel to the merger axis. We calculate the mean values in strips spaced every 100 kpc. Although the gas density and the magnetic field are peaked around the merger axis, the temperature and the turbulent Mach number are nearly constant. The 140 MHz intensity also exhibits a peaked profile, although it shows scatter around the peak. Fitting the profile with a Gaussian function, we obtain the best-fit full width half maximum (FWHM) of 0.98 Mpc. The bridge width depends on the projected direction. When the system is projected along the X axis (Fig. 15), the best-fit FWHM becomes 0.84 Mpc.
@@ 3.4 Point-point correlations
The pixel-by-pixel comparison of the radio and X-ray intensity maps provides important hints about the connection between thermal and non-thermal components in the ICM. Some classical radio halos in GCs are known to show sub-linear correlations, i.e., $I_{\rm radio}\propto I_{X}^{a_{\rm X}}$ with $a_{\rm X}<1$ (e.g., Govoni et al. 2001; Feretti et al. 2001; Bonafede et al. 2021). On the other hand, the correlation for mini halos tends to be super-linear ($a_{\rm X}>1$) (e.g., Ignesti et al. 2020; Biava et al. 2021). There are some exceptions, such as giant radio halos with super-linear correlations (e.g., Rajpurohit et al. 2018). The correlation in the radio bridge is still under debate due to the limited number of observations and data points available for each system. Analysing the LOFAR data of the A399–A401 bridge, de Jong et al. (2022) reported a weak sub-linear correlation between the radio and X-ray intensities with a slope of $a_{\rm X}=0.27\pm 0.07$ and a Spearman correlation coefficient of $r_{\rm S}=0.41\pm 0.14$. A similar correlation is also found for the bridge in A1758 ($a_{\rm X}=0.25$, $r_{\rm S}=0.52$, Botteon et al. 2020), which is also considered a pre-merger system. The point-to-point correlation between the radio intensity and the Compton-y parameter, which characterizes the thermal Sunyaev-Zeldovich (SZ) effect of ICM, has also been studied for a few nearby systems including A399–A401 (e.g., Planck Collaboration et al. 2013a; Radiconi et al. 2022). Hereafter, we call the correlation between the radio and X-ray intensities and that between the radio intensity and the Compton-y parameter as the radio–X correlation and the radio–SZ correlation, respectively.
In this section, we explore the radio-X and radio-SZ correlations in the simulated radio bridge, using 140 MHz and 3 keV as the reference frequency and photon energy for the radio and X-ray emission, respectively. For the X-ray emission, we ignore the line emission and only consider free-free emission for simplicity. We calculate the radio and X-ray intensities on a mesh with a resolution of 16 kpc, as in Fig. 4. The Compton-$y$ parameter is calculated as:
@@ 4.1 Evolution of a radio halo
Table 2: Additional models for the parameter study in Sect. 4.2. We fix $L=150$ kpc and $z_{i}=1$ in all of the models.
@@ 4.3 Limitations and caveats
In this exploratory work, we focus on the reacceleration of CREs in the filament region between a pair of merging clusters. Although a number of shocks are generated by the infall of mass clumps in the filament, the effect of shock (re)acceleration is not considered in this work. Govoni et al. (2019) discussed the shock acceleration scenario as a possible mechanism for the radio bridge and concluded that the reacceleration at weak shocks can explain the observed morphology under peculiar conditions (see Sect. 1). However, the efficiency of CR acceleration at such weak shocks is still under debate and may depend on the Alfvénic Mach number and the obliquity of the magnetic field (e.g., Kang et al. 2007; Bykov et al. 2019; Ryu et al. 2019; Gupta et al. 2025). Cosmological MHD simulations with spectrally-resolved CR modeling provides an opportunity to investigate the phenomenological constrains on the efficiency of shock acceleration in cluster environment (e.g., Böss et al. 2023; Vazza et al. 2023; Böss et al. 2025; Ivleva et al. 2026). Our tracer method onboard Enzo will be extended to include the on-the-fly shock finder to explore the CR acceleration at shocks.
@@ 5 Conclusions
An increasing number of observations report diffuse radio emission in cosmic filaments connecting merging clusters or groups. The mechanism responsible for accelerating CR particles in such low-density environment remains elusive. In this work, we extend previous studies using advanced numerical simulations to investigate the turbulent reacceleration of CREs in the inter-cluster region and the formation of a radio bridge. Using the AMR code Enzo, we conduct a cosmological MHD simulation of a system similar to A399–A401, which is known to host a 3 Mpc-long radio bridge. We upgrade the tracer method onboard Enzo and generate $N=1.5\times 10^{7}$ particles to sample the baryonic mass of $M_{\rm gas}=3\times 10^{14}$ M. We assume a population of CREs associated with each tracer particle and follow their spectral evolution with our parallel FP solver.
@@ Acknowledgements.
This work has received funding from the European High Performance Computing Joint Undertaking (JU) under grant agreement No 101093441, supported by contributions from multiple member states. The authors would like to thank all project partners and collaborators for their contributions and ongoing commitment to the success of the SPACE project. K.N. and G.B. acknowledge financial support from INAF grant ”Theory and simulations of non-thermal phenomena in galaxy clusters and beyond”. We acknowledge INAF for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy) and the LEONARDO consortium. The computations were conducted using the resources allocated to the projects designated ”INA24C5B12” and ”INA24C6B13”, with K.N. serving as the Principal Investigator. F.V. acknowledges the usage of online storage tools kindly provided by the INAF Astronomical Archive (IA2) initiative (http://www.ia2.inaf.it). Numerical computations were in part carried out on PC cluster at the Center for Computational Astrophysics, National Astronomical Observatory of Japan.
@@ Appendix B Tracer generation method
When tracer particles are generated, the code measures the baryon mass of each cell as $M_{\rm cell}=\rho_{\rm gas}(\Delta x/(1+z))^{3}$, where $\Delta x$ is the comoving cell width. The computed cell mass is then compared with $M_{\rm trc}^{\rm min}$ and $M_{\rm trc}^{\rm max}$. In low-resolution cells, the cell mass can exceed $M_{\rm trc}^{\rm max}$. In this case, we generate $\left\lfloor\frac{M_{\rm cell}}{M_{\rm trc}^{\rm max}}\right\rfloor$ tracers with mass $M_{\rm trc}=M_{\rm trc}^{\rm max}$, and, if the residual mass exceeds $M_{\rm trc}^{\rm min}$, an additional one tracer with the residual mass $M_{\rm trc}=M_{\rm cell}-M_{\rm trc}^{\rm max}*\left\lfloor\frac{M_{\rm cell}}{M_{\rm trc}^{\rm max}}\right\rfloor$. When the cell mass falls between $M_{\rm trc}^{\rm min}$ and $M_{\rm trc}^{\rm max}$, a single tracer with $M_{\rm trc}=M_{\rm cell}$ is generated. When $M_{\rm cell}<M_{\rm trc}^{\rm min}$, the tracer mass is too large to represent the cell. In this case, we adopt a Monte-Carlo sampling approach to include the contribution of such low-mass cells in a statistical sense: a tracer particle with $M_{\rm trc}=M_{\rm trc}^{\rm min}$ is generated with probability of $P=M_{\rm cell}/M_{\rm trc}^{\rm min}<1$. Although this method smooths the small-scale mass distribution below the mass resolution $M_{\rm trc}^{\rm min}$, it conserves the baryonic mass statistically. For each tracer, three random numbers in the range $[0,1)$ are generated to assign its initial $(x,y,z)$ position within the host cell. Tracer particles are treated as passive (virtual) particles and do not contribute to the calculation of the gravitational potential or the MHD evolution.
@@ Appendix E Projection along X axis
Fig. 15 shows the maps of 140 MHz radio emission and gas density projected along X axis of the simulation. The Mpc-sized radio bridge can also be seen in this projection at $z\lesssim 0.15$. As seen in the projection along Z axis (Fig. 1), there is a small halo moving in -X direction. This small halo overlaps with a part of the bridge region when the system is projected along X axis. On the other hand, Fig. 15 shows there are no significant substructures along Z axis. Since the merger axis is nearly aligned to the Y axis, we get qualitatively similar results for the one-dimensional profiles (Sect. 3.3). Fitting to the 140 MHz profile perpendicular to the bridge (almost along the Z axis) with a Gaussian function, we find the bridge width of 0.84 Mpc.

### 08 Apr 2026 v2

@@ 3.2.2 Spectral index map
We find that both the observed and simulated distributions have peaks around $\alpha_{60}^{140}\approx-1.5$. The simulated distribution is shifted to slightly larger $\alpha_{60}^{140}$ when the observational sensitivity is taken into account. The observed distribution is more extended toward the large-$\alpha_{60}^{140}$ (flat-spectrum) than the simulated one [^2] Note that the observational error is relatively large for the flat-spectrum points ($\alpha_{60}^{140}>-1$) compared to the steeper spectrum ones. See Fig. 2 of Pignataro et al. (2024a).. One of the possible causes for this lack of flat spectra would be our simplified assumption on the CR injection. Although we only consider an aged population of CREs injected at $z=1$, there should also be CREs freshly accelerated at shocks and AGNs at later times. The mixing of continuous injection and reacceleration is expected to generate different components with a broader spectral distribution and a stretched spectrum (e.g., Brunetti et al. 2017; Nishiwaki et al. 2021). In addition, our magnetic-field dynamo model may underestimate the dispersion of the magnetic field strength in the bridge region compared to the resolved dynamo in the MHD simulation (see Fig. 2). A larger field fluctuation may lead to a broader distribution of the spectral index.
@@ 3.3 Bridge profiles
The right panels of Fig. 8 show the profiles perpendicular to the projected merger axis. We consider a region with a width of 2 Mpc centered on the merger axis and a length of 1.2 Mpc parallel to the merger axis. We calculate the mean values in strips spaced every 100 kpc. Although the gas density and the magnetic field are peaked around the merger axis, the temperature and the turbulent Mach number are nearly constant. The 140 MHz intensity also exhibits a peaked profile, although it shows scatter around the peak. Fitting the profile with a Gaussian function, we obtain the best-fit full width at half maximum (FWHM) of 0.98 Mpc. The bridge width depends on the projected direction. When the system is projected along the X axis (Fig. 15), the best-fit FWHM becomes 0.84 Mpc.
@@ 3.4 Point-point correlations
The pixel-by-pixel comparison of the radio and X-ray intensity maps provides important hints about the connection between thermal and non-thermal components in the ICM. Some classical radio halos in GCs are known to show sub-linear correlations, i.e., $I_{\rm radio}\propto I_{X}^{a_{\rm X}}$ with $a_{\rm X}<1$ (e.g., Govoni et al. 2001; Feretti et al. 2001; Bonafede et al. 2021). On the other hand, the correlation for mini halos tends to be super-linear ($a_{\rm X}>1$) (e.g., Ignesti et al. 2020; Biava et al. 2021). There are some exceptions, such as giant radio halos with super-linear correlations (e.g., Rajpurohit et al. 2018). The correlation in the radio bridge is still under debate due to the limited number of observations and data points available for each system. Analyzing the LOFAR data of the A399–A401 bridge, de Jong et al. (2022) reported a weak sub-linear correlation between the radio and X-ray intensities with a slope of $a_{\rm X}=0.27\pm 0.07$ and a Spearman correlation coefficient of $r_{\rm S}=0.41\pm 0.14$. A similar correlation is also found for the bridge in A1758 ($a_{\rm X}=0.25$, $r_{\rm S}=0.52$, Botteon et al. 2020), which is also considered a pre-merger system. The point-to-point correlation between the radio intensity and the Compton-y parameter, which characterizes the thermal Sunyaev-Zeldovich (SZ) effect of ICM, has also been studied for a few nearby systems including A399–A401 (e.g., Planck Collaboration et al. 2013a; Radiconi et al. 2022). Hereafter, we call the correlation between the radio and X-ray intensities and that between the radio intensity and the Compton-y parameter as the radio–X correlation and the radio–SZ correlation, respectively.
In this section, we explore the radio-X and radio-SZ correlations in the simulated radio bridge, using 140 MHz and 3 keV as the reference frequency and photon energy for the radio and X-ray emissions, respectively. For the X-ray emission, we ignore the line emission and only consider free-free emission for simplicity. We calculate the radio and X-ray intensities on a mesh with a resolution of 16 kpc, as in Fig. 4. The Compton-$y$ parameter is calculated as:
@@ 4.1 Evolution of a radio halo
Table 2: Additional models for the parameter study in Sect. 4.2.333We fix $L=150$ kpc and $z_{i}=1$ in all of the models.
@@ 4.3 Limitations and caveats
In this exploratory work, we focus on the reacceleration of CREs in the filament region between a pair of merging clusters. Although a number of shocks are generated by the infall of mass clumps in the filament, the effect of shock (re)acceleration is not considered in this work. Govoni et al. (2019) discussed the shock acceleration scenario as a possible mechanism for the radio bridge and concluded that the reacceleration at weak shocks can explain the observed morphology under peculiar conditions (see Sect. 1). However, the efficiency of CR acceleration at such weak shocks is still under debate and may depend on the Alfvénic Mach number and the obliquity of the magnetic field (e.g., Kang et al. 2007; Bykov et al. 2019; Ryu et al. 2019; Gupta et al. 2025). Cosmological MHD simulations with spectrally-resolved CR modeling provides an opportunity to investigate the phenomenological constraints on the efficiency of shock acceleration in the cluster environment (e.g., Böss et al. 2023; Vazza et al. 2023; Böss et al. 2025; Ivleva et al. 2026). Our tracer method onboard Enzo will be extended to include the on-the-fly shock finder to explore the CR acceleration at shocks.
@@ 5 Conclusions
An increasing number of observations report diffuse radio emission in cosmic filaments connecting merging clusters or groups. The mechanism responsible for accelerating CR particles in such low-density environments remains elusive. In this work, we extend previous studies using advanced numerical simulations to investigate the turbulent reacceleration of CREs in the inter-cluster region and the formation of a radio bridge. Using the AMR code Enzo, we conduct a cosmological MHD simulation of a system similar to A399–A401, which is known to host a 3 Mpc-long radio bridge. We upgrade the tracer method onboard Enzo and generate $N=1.5\times 10^{7}$ particles to sample the baryonic mass of $M_{\rm gas}=3\times 10^{14}$ M. We assume a population of CREs associated with each tracer particle and follow their spectral evolution with our parallel FP solver.
@@ Acknowledgements.
We thank the anonymous referee for the careful review of the manuscript and for the constructive comments and suggestions, which helped improve the quality and clarity of this paper. This work has received funding from the European High Performance Computing Joint Undertaking (JU) under grant agreement No 101093441, supported by contributions from multiple member states. The authors would like to thank all project partners and collaborators for their contributions and ongoing commitment to the success of the SPACE project. K.N. and G.B. acknowledge financial support from INAF grant ”Theory and simulations of non-thermal phenomena in galaxy clusters and beyond”. F.V. acknowledges funding under the European Union’s Horizon Europe program through the ERC Synergy Grant COSMOMAG (Project Id. 101224803). We acknowledge INAF for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy) and the LEONARDO consortium. The computations were conducted using the resources allocated to the projects designated ”INA24C5B12” and ”INA24C6B13”, with K.N. serving as the Principal Investigator. F.V. acknowledges the usage of online storage tools kindly provided by the INAF Astronomical Archive (IA2) initiative (http://www.ia2.inaf.it). Numerical computations were in part carried out on PC cluster at the Center for Computational Astrophysics, National Astronomical Observatory of Japan.
@@ Appendix B Tracer generation method
When tracer particles are generated, the code measures the baryon mass of each cell as $M_{\rm cell}=\rho_{\rm gas}(\Delta x/(1+z))^{3}$, where $\Delta x$ is the comoving cell width. The computed cell mass is then compared with $M_{\rm trc}^{\rm min}$ and $M_{\rm trc}^{\rm max}$. In low-resolution cells, the cell mass can exceed $M_{\rm trc}^{\rm max}$. In this case, we generate $\left\lfloor\frac{M_{\rm cell}}{M_{\rm trc}^{\rm max}}\right\rfloor$ tracers with mass $M_{\rm trc}=M_{\rm trc}^{\rm max}$, and, if the residual mass exceeds $M_{\rm trc}^{\rm min}$, one additional tracer with the residual mass $M_{\rm trc}=M_{\rm cell}-M_{\rm trc}^{\rm max}*\left\lfloor\frac{M_{\rm cell}}{M_{\rm trc}^{\rm max}}\right\rfloor$. When the cell mass falls between $M_{\rm trc}^{\rm min}$ and $M_{\rm trc}^{\rm max}$, a single tracer with $M_{\rm trc}=M_{\rm cell}$ is generated. When $M_{\rm cell}<M_{\rm trc}^{\rm min}$, the tracer mass is too large to represent the cell. In this case, we adopt a Monte-Carlo sampling approach to include the contribution of such low-mass cells in a statistical sense: a tracer particle with $M_{\rm trc}=M_{\rm trc}^{\rm min}$ is generated with probability of $P=M_{\rm cell}/M_{\rm trc}^{\rm min}<1$. Although this method smooths the small-scale mass distribution below the mass resolution $M_{\rm trc}^{\rm min}$, it conserves the baryonic mass statistically. For each tracer, three random numbers in the range $[0,1)$ are generated to assign its initial $(x,y,z)$ position within the host cell. Tracer particles are treated as passive (virtual) particles and do not contribute to the calculation of the gravitational potential or the MHD evolution.
@@ Appendix E Projection along X axis
Fig. 15 shows the maps of 140 MHz radio emission and gas density projected along X axis of the simulation. The Mpc-sized radio bridge can also be seen in this projection at $z\lesssim 0.15$. As seen in the projection along Z axis (Fig. 1), there is a small halo moving in the -X direction. This small halo overlaps with a part of the bridge region when the system is projected along X axis. On the other hand, Fig. 15 shows that there are no significant substructures along the Z axis. Since the merger axis is nearly aligned to the Y axis, we get qualitatively similar results for the one-dimensional profiles (Sect. 3.3). Fitting to the 140 MHz profile perpendicular to the bridge (almost along the Z axis) with a Gaussian function, we find the bridge width of 0.84 Mpc.